OpenStudy (anonymous):

Hey im studying for a math test but im not sure if im ready yet. Help? which ratio forms a proportion? 4 over 12, 6 over 8 16 over 12, 3 over 4 3 over 2, 15 over 14 10 over 16, 5 over 8 can you help me and show me how?

OpenStudy (anonymous):

Wait, are these fractions?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Note that a proportion is a statement that two ratios are equivalent to each other. Two ratios form a proportion if their cross products are equal.

OpenStudy (anonymous):

Let's use for example, the following. \[\frac{ 3 }{ 5 } ~and~ \frac{ 6 }{ 10 }\]

OpenStudy (anonymous):

To solve that, you want to compare the two cross products. Multiply the numerator of one fraction and the denominator of the other. So it will be setup like this: 3 × 10 = 5 × 6, is that true or false? Well, 3 × 10 = 30 and 5 × 6 = 30 So 30 = 30. The question was "Do these ratios form a proportion?" the correct answer for this example is, yes. You see now? :)

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

Okay, so let's set it up like so... \[\frac{ 4 }{ 12 }~and~\frac{ 6 }{ 8 }\] Now what would we multiply to solve that?

OpenStudy (anonymous):

Okay Well, I got to go. I hope I assisted you.